OpenStudy (anonymous):

Help me simplify? x-4/2x-6 + 3/x^2-9

OpenStudy (anonymous):

could you find the common factor between two fraction ?

OpenStudy (zzr0ck3r):

2x-6 = 2(x-3) x^2-9 = (x-3)(x+3)

OpenStudy (anonymous):

they have x-3 in common

OpenStudy (zzr0ck3r):

your common denominator should be 2(x-3)(x+3)

OpenStudy (anonymous):

So my solution would be x-2/2(x-3)?